OpenStudy (anonymous):

COuld someone help me with this... Which of the following is a solution of x2 + 6x = -22?

hartnn (hartnn):

do you know the quadratic formula ?

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\) did you follow these steps ?

hartnn (hartnn):

a=1,b=6,c=22

hartnn (hartnn):

b^2-4ac = 36-4*22 = 36 -88= -52 oh, you got your error ? try again and see whether you get it now :)

hartnn (hartnn):

what about the denominator 2a = 2 ?

hartnn (hartnn):

thats right, none of the choices are correct

hartnn (hartnn):

-6 +/- sqrt (-52) / 2 = -3 +/- i sqrt 13 could you get how i got that final answer ?

hartnn (hartnn):

thats the correct answer and its not in choices.... so either the question is incorrect or the options...

hartnn (hartnn):

i did use the imaginary number, \(\sqrt{-52}= \sqrt{-1}\sqrt{52}=i \sqrt {52}=2i \sqrt 13\)